# 22mm22mm25mm25mm

### v1

@@ 2.4. Algorithm for simultaneous block diagonalization
3:Caculate the center $Z(A_{1},A_{2},\ldots,A_{m})$ by solving system $(A_{i}X)^{T}=A_{i}X$ for $1\leq i\leq m$. Select base matrices $X_{j}$ for $1\leq j\leq\operatorname{dim}{Z(A_{1},A_{2},\ldots,A_{m})}$.
@@
4:Compute a pair of orthogonal idempotents $\epsilon_{1}$ and $\epsilon_{2}$ using $Z(A_{1},A_{2},\ldots,A_{m})$. If they do not exist, $A_{1},A_{2},\ldots,A_{m}$ are not simultaneously block diagonalizable via congruence. Otherwise, continue.
@@
5:Compute invertible matrix $P$ such that $P^{-1}\epsilon_{1}P=\begin{pmatrix}I_{1}&\\ &0\end{pmatrix}$ and $P^{-1}\epsilon_{2}P=\begin{pmatrix}0&\\ &I_{2}\end{pmatrix}$. Then simultaneously block diagonalize $A_{1},A_{2},\ldots,A_{m}$ such that $P^{T}A_{i}P$ takes the form of a block diagonal matrix $\begin{pmatrix}B_{i1}&\\ &B_{i2}\end{pmatrix}$ for $1\leq i\leq m$. Go to step 1 and continue to simultaneously block diagonalize $B_{11},B_{21},\ldots,B_{m1}$ and $B_{12},B_{22},\ldots,B_{m2}$ via congruence separately.
@@
Since these submatrices may be simultaneously block diagonalizable via congruence, Algorithm 1 is called again using the submatrices as input. This process occurs recursively until all output matrices are not simultaneously block diagonalizable via congruence. The resulting output matrices can then be reassembled as block diagonal matrices corresponding to the original symmetric matrices $A_{1},A_{2},\ldots,A_{m}$. During this process, the maximum number of blocks for the SBDC is also determined.
@@
This algorithm contains several components, including computations of centers and orthogonal idempotents, which use standard tasks in linear algebra [14]. Calculating the centers of symmetric matrices requires solving a system of linear equations, which has polynomial complexity [5]. The orthogonal idempotents are obtained by imposing constraints on the rank and trace of matrix $X$ in Equation (2.1), which leads to a system of nonlinear equations. For the SDC of a set of symmetric matrices, this system only contains sparse quadratic equations that are solved efficiently in numerical experiments. We refer to [5] for more information. The complexity analysis of the SBDC of a set of symmetric matrices is still unavailable and will be addressed in future studies.
3. More examples
We provide more examples to demonstrate the procedure of this technique. The SDC of two and three symmetric matrices are shown in Examples 3.1 and 3.2, respectively. The SBDC of two and three symmetric matrices are given in Examples 3.3 and 3.4, respectively.
@@ 3. More examples
(Example 2.3 continued) Consider symmetric matrices
@@
Using the center calculated in Example 2.3, we obtain base matrices of $Z(A_{1},A_{2})$ as
$X_{1}=\begin{pmatrix}1&-1&1\\ 6&6&0\\ 0&0&0\end{pmatrix},\quad X_{2}=\begin{pmatrix}-5&-1&-2\\ 6&0&3\\ 0&0&0\end{pmatrix},\quad X_{3}=\begin{pmatrix}-1&0&0\\ 1&0&0\\ 2&0&0\end{pmatrix},$
$X_{4}=\begin{pmatrix}-1&-5&-1\\ -6&0&0\\ 0&12&0\end{pmatrix},\quad X_{5}=\begin{pmatrix}5&1&-2\\ -6&0&0\\ 0&0&6\end{pmatrix}.$
Since $-X_{3}$ is a primitive idempotent, we set $\epsilon_{1}=-X_{3}$ and $\epsilon_{2}=I-\epsilon_{1}$. A pair of orthogonal idempotents is produced as
@@
$A_{1}$ and $A_{2}$ can then be simultaneously block diagonalized via congruence as
@@
respectively. Since neither matrices are fully diagonalized, we continue to simultaneously diagonalize their bottom-right submatrices, which are denoted as $B_{1}=\begin{pmatrix}2&0\\ 0&-1\end{pmatrix}$ and $B_{2}=\begin{pmatrix}5&1\\ 1&-1\end{pmatrix}$. Their center is calculated as
$Z(B_{1},B_{2})=\left\{\begin{pmatrix}-3a+b&-a\\ 2a&b\end{pmatrix}\vline\,a,b\in\mathbbm{k}\right\}.$
@@
Note that $1,1$-th entries of both transformed matrices are zero. This implies that $A_{1}$ and $A_{2}$ are degenerate matrices.
@@
$A_{1}=\begin{pmatrix}1&1&1\\ 1&3&3\\ 1&3&15\end{pmatrix},\quad A_{2}=\begin{pmatrix}1&1&1\\ 1&4&4\\ 1&4&20\end{pmatrix},\quad A_{3}=\begin{pmatrix}1&1&1\\ 1&5&5\\ 1&5&25\end{pmatrix}.$
We compute their center by Equation (2.1) as
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}a+b+c&a&a\\ 0&b+c&b\\ 0&0&c\end{pmatrix}\vline\,a,b,c\in\mathbbm{k}\right\}$
@@
$X_{1}=\begin{pmatrix}1&1&1\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad X_{2}=\begin{pmatrix}1&0&0\\ 0&1&1\\ 0&0&0\end{pmatrix},\quad X_{3}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
@@
$\epsilon_{1}=\begin{pmatrix}1&1&1\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}0&-1&-1\\ 0&1&0\\ 0&0&1\end{pmatrix}$
@@
where $P_{1}=\begin{pmatrix}1&-1&-1\\ 0&1&0\\ 0&0&1\end{pmatrix}$. Thus, we can simultaneously block diagonalize the symmetric matrices via congruence as
$P_{1}^{T}A_{1}P_{1}=\begin{pmatrix}1&&\\ &2&2\\ &2&14\end{pmatrix},\>\quad P_{1}^{T}A_{2}P_{1}=\begin{pmatrix}1&&\\ &3&3\\ &3&19\end{pmatrix},\>\quad P_{1}^{T}A_{3}P_{1}=\begin{pmatrix}1&&\\ &4&4\\ &4&24\end{pmatrix}.$
We repeat this process using their submatrices, which are denoted as
$B_{1}=\begin{pmatrix}2&2\\ 2&14\end{pmatrix},\>\quad B_{2}=\begin{pmatrix}3&3\\ 3&19\end{pmatrix},\>\quad B_{3}=\begin{pmatrix}4&4\\ 4&24\end{pmatrix}.$
Their center is calculated as
$Z(B_{1},B_{2},B_{3})=\left\{\begin{pmatrix}a+b&a\\ 0&b\\ \end{pmatrix}\vline\,a,b\in\mathbbm{k}\right\}.$
We obtain a new pair of orthogonal idempotent $\epsilon_{3}=\begin{pmatrix}1&1\\ 0&0\end{pmatrix}$ and $\epsilon_{4}=\begin{pmatrix}0&-1\\ 0&1\end{pmatrix}$, then derive a corresponding invertible matrix $P_{2}=\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}$ such that $P_{2}^{T}\epsilon_{3}P_{2}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ and $P_{2}^{T}\epsilon_{4}P_{2}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$. Thus, $B_{1},B_{2},B_{3}$ are simultaneously diagonalized via congruence as
$P_{2}^{T}B_{1}P_{2}=\begin{pmatrix}2&0\\ 0&12\end{pmatrix},\>\quad P_{2}^{T}B_{2}P_{2}=\begin{pmatrix}3&0\\ 0&16\end{pmatrix},\>\quad P_{2}^{T}B_{3}P_{2}=\begin{pmatrix}4&0\\ 0&20\end{pmatrix},$
respectively. Finally, $A_{1},A_{2},A_{3}$ are simultaneously diagonalized via congruence as
$P^{T}A_{1}P=\begin{pmatrix}1&&\\ &2&\\ &&12\end{pmatrix},\>\quad P^{T}A_{2}P=\begin{pmatrix}1&&\\ &3&\\ &&16\end{pmatrix},\>\quad P^{T}A_{3}P=\begin{pmatrix}1&&\\ &4&\\ &&20\end{pmatrix},$
@@
$P=P_{1}\begin{pmatrix}1&0\\ 0&P_{2}\end{pmatrix}=\begin{pmatrix}1&0&0\\ 0&2&0\\ 0&2&12\end{pmatrix}.$
@@
Let symmetric matrices be
$A_{1}=\begin{pmatrix}1&1&0\\ 1&3&6\\ 0&6&16\end{pmatrix},\quad A_{2}=\begin{pmatrix}1&1&0\\ 1&4&8\\ 0&8&20\end{pmatrix}.$
We calculate their center according to Equation (2.1) as
$Z(A_{1},A_{2})=\left\{\begin{pmatrix}a+2b-4c+18d&a&4c-51d\\ -2b&2b-4c+18d&4b-4c+51d\\ b&c&d\end{pmatrix}\vline\,a,b,c,d\in\mathbbm{k}\right\}$
and work out a set of base matrices
$X_{1}=\begin{pmatrix}1&1&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\enspace X_{2}=\begin{pmatrix}2&0&0\\ -2&2&4\\ 1&0&0\end{pmatrix},\enspace X_{3}=\begin{pmatrix}-4&0&4\\ 0&-4&-4\\ 0&1&0\end{pmatrix},\enspace X_{4}=\begin{pmatrix}18&0&-51\\ 0&18&51\\ 0&0&1\end{pmatrix}.$
Based on the algebraic structure of $Z(A_{1},A_{2})$, we compute orthogonal idempotents
$\epsilon_{1}=\begin{pmatrix}1&1&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}0&-1&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$
and
$P^{-1}\epsilon_{1}P=\begin{pmatrix}1&&\\ &0&\\ &&0\end{pmatrix},\quad P^{-1}\epsilon_{2}P=\begin{pmatrix}0&&\\ &1&\\ &&1\end{pmatrix},$
where $P=\begin{pmatrix}1&-1&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$.
By simultaneously block diagonalizing $A_{1}$ and $A_{2}$ via congruence using $P$, we obtain
$P^{T}A_{1}P=\begin{pmatrix}1&&\\ &3&8\\ &8&20\end{pmatrix},\quad P^{T}A_{2}P=\begin{pmatrix}1&&\\ &2&6\\ &6&16\end{pmatrix},$
which may be further simultaneously diagonalized via congruence. We denote their bottom-right submatrices as
$B_{1}=\begin{pmatrix}3&8\\ 8&20\end{pmatrix},\quad B_{2}=\begin{pmatrix}2&6\\ 6&16\end{pmatrix},$
respectively, and compute their center as
$Z(B_{1},B_{2})=\left\{\begin{pmatrix}a&-4b\\ b&a+4b\end{pmatrix}\vline\,a,b\in\mathbbm{k}\right\}.$
Since $Z(B_{1},B_{2})\cong\mathbbm{k}[\varepsilon]/(\varepsilon^{2})$ does not have a nontrivial idempotent, $B_{1},B_{2}$ are not simultaneously diagonalizable via congruence. Thus, $A_{1},A_{2}$ are simultaneously block diagonalizable via congruence.
Example 3.4.
Suppose we have symmetric matrices
@@
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}a-8b+c&2a-b&3a\\ 0&-12b+c&-12b\\ 0&3b&c\end{pmatrix}\vline\,a,b,c\in\mathbbm{k}\right\}$
@@
$X_{1}=\begin{pmatrix}1&2&3\\ 0&0&0\\ 0&0&0\end{pmatrix},\enspace X_{2}=\begin{pmatrix}-8&-1&0\\ 0&-12&-12\\ 0&3&0\end{pmatrix},\enspace X_{3}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}.$
@@
$Z(B_{1},B_{2},B_{3})=\left\{\begin{pmatrix}4a+b&4a\\ -a&b\end{pmatrix}\vline\,a,b\in\mathbbm{k}\right\}$
@@
Since a nontrivial idempotent does not exist in $Z(B_{1},B_{2},B_{3})\cong\mathbbm{k}[\varepsilon]/(\varepsilon^{2})$, $B_{1},B_{2},B_{3}$ are not simultaneously diagonalizable via congruence and $A_{1},A_{2},A_{3}$ are simultaneously block diagonalizable via congruence as shown above.
4. Simultaneous block diagonalization of Hermitian matrices
In this section, we extend our results on the SBDC of a set of symmetric matrices to the SBDC of a set of Hermitian matrices. For the convenience of readers, we recall the definition of Hermitian matrices, see [12] for more details.

### v2

@@ 2.3 Algorithm for simultaneous block diagonalization
3:Calculate the center $Z(A_{1},A_{2},\ldots,A_{m})$ by solving system $(A_{i}X)^{T}=A_{i}X$ for $1\leq i\leq m$. Select base matrices $X_{j}$ for $1\leq j\leq\operatorname{dim}{Z(A_{1},A_{2},\ldots,A_{m})}$ such that $X=\sum_{j=1}^{\operatorname{dim}{Z(A_{1},A_{2},\ldots,A_{m})}}X_{j}$.
@@
4:Impose the constraint $X^{2}=X$ on the center $Z(A_{1},A_{2},\ldots,A_{m})$ to compute a nontrivial idempotent $\epsilon_{1}$. For orthogonal congruence, an additional constraint $X^{T}=X$ is imposed on $Z(A_{1},A_{2},\ldots,A_{m})$. If $\epsilon_{1}$ does not exist, $A_{1},A_{2},\ldots,A_{m}$ are not simultaneously block diagonalizable via congruence. Otherwise, calculate the orthogonal idempotent $\epsilon_{2}=I-\epsilon_{1}$ and continue.
5:Compute invertible matrix $P$ whose columns form a basis of simultaneous eigenvectors for $\epsilon_{1}$ and $\epsilon_{2}$, which guarantees $P^{-1}\epsilon_{1}P=\begin{pmatrix}I_{1}&\\ &0\end{pmatrix}$ and $P^{-1}\epsilon_{2}P=\begin{pmatrix}0&\\ &I_{2}\end{pmatrix}$. For orthogonal congruence, take bases of column spaces of $\epsilon_{1}$ and $\epsilon_{2}$, then apply Gram–Schmidt orthogonalization to obtain an orthogonal matrix $P$.
@@
6:Simultaneously block diagonalize $A_{1},A_{2},\ldots,A_{m}$ via congruence such that $P^{T}A_{i}P$ takes the form of a block diagonal matrix $\begin{pmatrix}B_{i1}&\\ &B_{i2}\end{pmatrix}$ for $1\leq i\leq m$. Go to step 1 and continue to simultaneously block diagonalize $B_{11},B_{21},\ldots,B_{m1}$ and $B_{12},B_{22},\ldots,B_{m2}$ via congruence separately.
@@
Since these submatrices may admit further SBDC, Algorithm 1 is applied recursively. This process occurs recursively until all output matrices are not simultaneously block diagonalizable via congruence. The resulting output matrices can then be reassembled as block diagonal matrices corresponding to the original symmetric matrices $A_{1},A_{2},\ldots,A_{m}$. The maximum number of blocks for this decomposition is uniquely determined by the structure of the primitive orthogonal idempotent elements in the associated Jordan algebra [mccrimmon2003taste].
@@
This algorithm contains several components, including computations of centers and orthogonal idempotents, which use standard tasks in linear algebra [huang2021centres]. Calculating the centers of symmetric matrices requires solving a system of linear equations, which has polynomial complexity [horn2012matrix]. The orthogonal idempotents are obtained by imposing constraints on the rank and trace of the matrix $X$ in Equation (2.1), which yields a system of nonlinear equations. For the SDC of a set of symmetric matrices, this system only contains sparse quadratic equations that are solved efficiently in numerical experiments [fang2025numerical]. For the general SBDC, determining idempotents may involve solving polynomial systems of higher complexity and remains a challenging problem. Nevertheless, for many practical instances with lower-dimensional centers, it remains computationally feasible.
3 More examples
This section presents additional examples to illustrate the procedure of this technique. While the theory is developed for a general field $\mathbbm{k}$, the following computations are performed over either $\mathbbm{R}$ or $\mathbbm{C}$ for the convenience of the exposition. The SDC of two and three symmetric matrices are shown in Examples 3.1 and 3.2, respectively. The SBDC and orthogonal SBDC of three symmetric matrices are given in Examples 3.3 and 3.4, respectively.
@@ 3 More examples
(Example 2.3 continued) Consider the following two real symmetric matrices
@@
Using the center calculated in Example 2.3, we obtain the base matrices of $Z(A_{1},A_{2})$ as
$X_{1}=\begin{pmatrix}1&-1&1\\ 6&6&0\\ 0&0&0\end{pmatrix},\,X_{2}=\begin{pmatrix}-5&-1&-2\\ 6&0&3\\ 0&0&0\end{pmatrix},\,X_{3}=\begin{pmatrix}-1&0&0\\ 1&0&0\\ 2&0&0\end{pmatrix},\,X_{4}=\begin{pmatrix}-1&-5&-1\\ -6&0&0\\ 0&12&0\end{pmatrix},\,X_{5}=\begin{pmatrix}5&1&-2\\ -6&0&0\\ 0&0&6\end{pmatrix}$.
For orthogonal congruence, we impose an additional symmetry condition $X^{T}=X$ on the idempotent $X\in Z(A_{1},A_{2})$. Since the only solutions are 0 and $I_{3}$, $A_{1}$ and $A_{2}$ are not simultaneously block diagonalizable via orthogonal congruence. Thus, we proceed to calculate orthogonal idempotents for general congruence. Since $X_{3}$ is a primitive idempotent, we set $\epsilon_{1}=-X_{3}$ and $\epsilon_{2}=I-\epsilon_{1}$. A pair of orthogonal idempotents is produced as
@@
The matrices $A_{1}$ and $A_{2}$ can then be simultaneously block diagonalized via congruence as
@@
respectively. Since neither matrix is fully diagonalized, we continue to simultaneously diagonalize their bottom-right submatrices, which are denoted as $B_{1}=\begin{pmatrix}2&0\\ 0&-1\end{pmatrix}$ and $B_{2}=\begin{pmatrix}5&1\\ 1&-1\end{pmatrix}$. Their center is calculated as
$Z(B_{1},B_{2})=\left\{\begin{pmatrix}-3a+b&-a\\ 2a&b\end{pmatrix}\vline\,a,b\in\mathbbm{R}\right\}.$
@@
It is interesting to note that $A_{1}$ and $A_{2}$ are not simultaneously decomposable via orthogonal congruence, but they are simultaneously diagonalizable via general congruence.
@@
$A_{1}=\begin{pmatrix}1&0&0\\ 0&-1&1\\ 0&1&2\end{pmatrix},\quad A_{2}=\begin{pmatrix}0&1&-1\\ 1&-1&1\\ -1&1&1\end{pmatrix},\quad A_{3}=\begin{pmatrix}2&1&-1\\ 1&-3&3\\ -1&3&-2\end{pmatrix}$.
Firstly, we consider their SBDC over $\mathbbm{R}$. We compute their center by Equation (2.1) as
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}a-b+c&-c&c\\ c&a-b&b\\ 0&0&a\end{pmatrix}\vline\,a,b\in\mathbbm{R}\right\}$
@@
$X_{1}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},\quad X_{2}=\begin{pmatrix}-1&0&0\\ 0&-1&1\\ 0&0&0\end{pmatrix},\quad X_{3}=\begin{pmatrix}1&-1&1\\ 1&0&0\\ 0&0&0\end{pmatrix}$.
@@
$\epsilon_{1}=\begin{pmatrix}1&0&0\\ 0&1&-1\\ 0&0&0\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&0&1\end{pmatrix}$
@@
$P_{1}^{-1}\epsilon_{1}P_{1}=\begin{pmatrix}1&&\\ &1&\\ &&0\end{pmatrix},\quad P_{1}^{-1}\epsilon_{2}P_{1}=\begin{pmatrix}0&&\\ &0&\\ &&1\end{pmatrix}$,
where $P_{1}=\begin{pmatrix}1&1&0\\ 0&1&1\\ 0&0&1\end{pmatrix}$. Thus, we can simultaneously block diagonalize the symmetric matrices via congruence as
$P_{1}^{T}A_{1}P_{1}=\begin{pmatrix}1&1&\\ 1&0&\\ &&3\end{pmatrix},\quad P_{1}^{T}A_{2}P_{1}=\begin{pmatrix}0&1&\\ 1&1&\\ &&2\end{pmatrix},\quad P_{1}^{T}A_{3}P_{1}=\begin{pmatrix}2&3&\\ 3&1&\\ &&1\end{pmatrix}$.
We repeat this process using their top left submatrices, which are denoted as
$B_{1}=\begin{pmatrix}1&1\\ 1&0\end{pmatrix},\quad B_{2}=\begin{pmatrix}0&1\\ 1&1\end{pmatrix},\quad B_{3}=\begin{pmatrix}2&3\\ 3&1\end{pmatrix}$.
Their center is computed as
$Z(B_{1},B_{2},B_{3})=\left\{\begin{pmatrix}a-b&-b\\ b&a\end{pmatrix}\vline\,a,b\in\mathbbm{R}\right\}.$
We obverse that $Z(B_{1},B_{2},B_{3})\cong\mathbbm{C}$ as algebras over $\mathbbm{R}$, in other words, it only contains trivial idempotent elements 0 and $I_{3}$. This is also straightforward by computation. Therefore, $B_{1},B_{2},B_{3}$ are not simultaneously diagonalizable via congruence over $\mathbbm{R}$ and $A_{1},A_{2},A_{3}$ are simultaneously block diagonalizable via congruence over $\mathbbm{R}$ as above.
Now we consider the SBDC problem over $\mathbbm{C}$, then $Z(A_{1},A_{2},A_{3})$ shares the same bases and can be simultaneously block diagonalized as above. In contrast, $Z(B_{1},B_{2},B_{3})\cong\mathbbm{C}\times\mathbbm{C}$ over the complex field $\mathbbm{C}$ and must contain nontrivial idempotents. We obtain a new pair of orthogonal idempotents $\epsilon_{3}=\begin{pmatrix}\frac{3-i\sqrt{3}}{6}&-\frac{i\sqrt{3}}{3}\\ \frac{i\sqrt{3}}{3}&\frac{3+i\sqrt{3}}{6}\end{pmatrix}$ and $\epsilon_{4}=\begin{pmatrix}\frac{3+i\sqrt{3}}{6}&\frac{i\sqrt{3}}{3}\\ -\frac{i\sqrt{3}}{3}&\frac{3-i\sqrt{3}}{6}\end{pmatrix}$, then derive a corresponding invertible matrix $P_{2}=\begin{pmatrix}\frac{-1-i\sqrt{3}}{2}&\frac{-1+i\sqrt{3}}{2}\\ 1&1\end{pmatrix}$ such that $P_{2}^{-1}\epsilon_{3}P_{2}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ and $P_{2}^{-1}\epsilon_{4}P_{2}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$. Thus, $B_{1},B_{2},B_{3}$ are simultaneously diagonalized via congruence as
$P_{2}^{T}B_{1}P_{2}=\begin{pmatrix}\frac{-3-i\sqrt{3}}{2}&\\ &\frac{-3+i\sqrt{3}}{2}\end{pmatrix},\>\quad P_{2}^{T}B_{2}P_{2}=\begin{pmatrix}-i\sqrt{3}&\\ &i\sqrt{3}\end{pmatrix},\>\quad P_{2}^{T}B_{3}P_{2}=\begin{pmatrix}-3-2i\sqrt{3}&\\ &-3+2i\sqrt{3}\end{pmatrix}$,
respectively. Finally, $A_{1},A_{2},A_{3}$ are simultaneously diagonalized via congruence over $\mathbbm{C}$ as
$P^{T}A_{1}P=\begin{pmatrix}\frac{-3-i\sqrt{3}}{2}&\\ &\frac{-3+i\sqrt{3}}{2}&\\ &&3\end{pmatrix},\quad P^{T}A_{2}P=\begin{pmatrix}-i\sqrt{3}&&\\ &i\sqrt{3}&\\ &&2\end{pmatrix},\quad P^{T}A_{3}P=\begin{pmatrix}-3-2i\sqrt{3}&&\\ &-3+2i\sqrt{3}&\\ &&1\end{pmatrix}$,
@@
$P=P_{1}\begin{pmatrix}P_{2}&0\\ 0&1\end{pmatrix}=\begin{pmatrix}\frac{1-i\sqrt{3}}{2}&\frac{1+i\sqrt{3}}{2}&0\\ 1&1&1\\ 0&0&1\end{pmatrix}$.
This example indicates that the SBDC problem depends on the ground field.
@@
Suppose we have three complex symmetric matrices
@@
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}a-8b+c&2a-b&3a\\ 0&-12b+c&-12b\\ 0&3b&c\end{pmatrix}\vline\,a,b,c\in\mathbbm{C}\right\}$
@@
$X_{1}=\begin{pmatrix}1&2&3\\ 0&0&0\\ 0&0&0\end{pmatrix},\kern 5.0ptX_{2}=\begin{pmatrix}-8&-1&0\\ 0&-12&-12\\ 0&3&0\end{pmatrix},\kern 5.0ptX_{3}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix}$.
@@
$Z(B_{1},B_{2},B_{3})=\left\{\begin{pmatrix}4a+b&4a\\ -a&b\end{pmatrix}\vline\,a,b\in\mathbbm{C}\right\}$
@@
One can verify by direct computation that there are no nontrivial idempotents in $Z(B_{1},B_{2},B_{3})$. This also follows from the center’s algebraic structure $Z(B_{1},B_{2},B_{3})\cong\mathbbm{C}[\varepsilon]/(\varepsilon^{2})$. Thus, $B_{1},B_{2},B_{3}$ are not simultaneously diagonalizable via congruence, and $A_{1},A_{2},A_{3}$ are simultaneously block diagonalizable via congruence as shown above.
Example 3.4.
Let real symmetric matrices be
$A_{1}=\begin{pmatrix}-9&4&12\\ 4&10&3\\ 12&3&-16\end{pmatrix},\quad A_{2}=\begin{pmatrix}16&8&12\\ 8&5&6\\ 12&6&9\end{pmatrix},\quad A_{3}=\begin{pmatrix}41&-4&12\\ -4&20&-3\\ 12&-3&34\end{pmatrix}$.
We compute their center according to Equation (2.1) as
$Z(A_{1},A_{2},A_{3})=\left\{\begin{pmatrix}a+9b&0&-12b\\ 0&a&0\\ -12b&0&a+16b\end{pmatrix}\vline\,a,b\in\mathbbm{R}\right\}$
and work out two base matrices
$X_{1}=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},\quad X_{2}=\begin{pmatrix}9&0&-12\\ 0&0&0\\ -12&0&16\end{pmatrix}$.
To perform the orthogonal SBDC, we impose the symmetric condition on a non-trivial idempotent element from $Z(A_{1},A_{2},A_{3})$. Since $X_{2}^{2}=25X_{2}$, we normalize $X_{2}$ and obtain a pair of orthogonal idempotents
$\epsilon_{1}=\begin{pmatrix}\frac{9}{25}&0&-\frac{12}{25}\\ 0&0&0\\ -\frac{12}{25}&0&\frac{16}{25}\end{pmatrix},\quad\epsilon_{2}=\begin{pmatrix}\frac{16}{25}&0&\frac{12}{25}\\ 0&1&0\\ \frac{12}{25}&0&\frac{9}{25}\end{pmatrix}$.
Thus, we compute an orthogonal matrix $P=\frac{1}{5}\begin{pmatrix}3&0&4\\ 0&5&0\\ -4&0&3\end{pmatrix}$ such that
By simultaneously block diagonalizing $A_{1},A_{2},A_{3}$ via orthogonal congruence, we obtain
$P^{T}A_{1}P=\begin{pmatrix}-25&&\\ &10&5\\ &5&0\end{pmatrix},\quad P^{T}A_{2}P=\begin{pmatrix}0&&\\ &5&10\\ &10&25\end{pmatrix},\quad P^{T}A_{3}P=\begin{pmatrix}25&&\\ &20&-5\\ &-5&50\end{pmatrix}$,
which may be further simultaneously diagonalized via orthogonal congruence. We denote their bottom-right submatrices as
$B_{1}=\begin{pmatrix}10&5\\ 5&0\end{pmatrix},\quad B_{2}=\begin{pmatrix}5&10\\ 10&25\end{pmatrix},\quad B_{2}=\begin{pmatrix}20&-5\\ -5&50\end{pmatrix}$,
respectively, and compute their center as
$Z(B_{1},B_{2},B_{3})=\left\{\begin{pmatrix}a&0\\ 0&a\end{pmatrix}\vline\,a\in\mathbbm{R}\right\}.$
Since a nontrivial idempotent does not exist in $Z(B_{1},B_{2},B_{3})\cong\mathbbm{R}$, they are not simultaneously diagonalizable via congruence, and $A_{1},A_{2},A_{3}$ are simultaneously block diagonalizable via orthogonal congruence as shown above.
4 Simultaneous block diagonalization of Hermitian matrices
In this section, we extend our results on the SBDC of a set of symmetric matrices to the SBDC of a set of Hermitian matrices. We briefly recall the definition of Hermitian matrices, see [horn2012matrix] for more details.